[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Oct 24, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Oct 24 14:06:12 UTC 2017

Here is today's summary of NHC computer operations:
--- Matt Onderlinde


1) AWIPS Security Patch from 930-1035am. Jeff disconnected from  
collaboration server. He tried to log back in, but was unable (login  
froze). "Unable to get bookmarked rooms from server" alertviz  
appeared, "Problem registering Public Key" also.

2) 12z Pacific NWPS only came in for initial timestep despite run  
completing out to 168 hours. Mello investigating and checking to see if  
this is related to the security patch this morning.
ECMWF was approximately 3 hours late at 00Z (Oct 24). This issue was with  
ECMWF, no TSB action is required.
Other items:
1) Building AC System turned on with Fans running at highest level in all  
ops areas but warm/hot air blowing continuously. JL notified William at  
1215pm. William was unaware of the issue but said that he would look into  
it. So far the server room is OK with 1 cooling unit at 69F, another at  
70F, and a third slightly warm at 74F.
No RFCs scheduled for implementation this week, see NHC RFC Memo 2017 10  
20.pdf for NCO RFCs that may affect NHC operations.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171024/87993e44/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list